Find the length of the hypotenuse of a right triangle with legs of lengths 6 and 8.
Answer by mananth(16946)

one leg = 6
other leg = 8
so 6^2+8^2= hyp^2
36+64 = Hyp^2
100= Hyp^2
take the square root
hypotenuse = 10

...
m.ananth@hotmail.ca